### v1

# Daejeon 34141, South Korea

Abstract.
Let $N$ be a manifold of dimension $m$ with a flat vector bundle given by a representation $\rho:\pi_{1}(N)\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ where $\pi_{1}(N)$ is finitely generated. The holonomy group $\rho$ is a $k$-partially hyperbolic holonomy representation if the flat bundle pulled back over the unit tangent bundle of a sufficiently large compact submanifold of $N$ splits into expanding, neutral, and contracting subbundles along the geodesic flow, where the expanding and contracting subbundles are $k$-dimensional with $k<n/2$. Suppose that each element of $\rho(\pi_{1}(N))$ has an eigenvalue of norm $1$, or, alternatively, each element of it has a singular value that is uniformly bounded above and below. We show that $\rho$ is a $P$-Anosov representation for a parabolic subgroup $P$ of ${\mathsf{GL}}(n,{\mathbb{R}})$ if and only if $\rho$ is a partially hyperbolic representation. We are going to primarily employ representation theory techniques. As an application, we will show this holds when $N$ is a complete affine $n$-manifold and $\rho$ is a linear part of the holonomy representation. This had never been done over the full general linear group.
Key words and phrases:
affine manifold, Anosov representation, word-hyperbolic group, Gromov hyperbolic space, coarse geometry
2010 Mathematics Subject Classification:
Primary 57M50; Secondary 22E40
This research was supported by the Basic Science Research Program through the National Research Foundation of Korea(NRF) funded by the Ministry of Education(2019R1A2C108454412), Author orcid number: 0000-0002-9201-8210
1. Introduction
1.1. Notation
Let $N$ be an $m$-manifold with a finitely generated fundamental group $\pi_{1}(N)$. Each flat vector bundle over $N$ can be described as ${\mathbb{R}}^{n}_{\rho}(N):=\tilde{N}\times{\mathbb{R}}^{n}/\sim$ where $(x,\vec{v})\sim(\gamma(x),\rho(\gamma)(\vec{v})$ for $\gamma\in\pi_{1}(N)$ where $\rho:\pi_{1}(N)\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ is a representation. A differential-geometry description of a flat bundle is a vector bundle with a flat linear connection.
Let $M$ be a compact submanifold (with or without boundary) in $N$. We assume that the inclusion map $M\hookrightarrow N$ induces a surjection on fundamental groups. The inverse image $\hat{M}$ of $M$ in $\tilde{N}$ is connected, and the deck transformation group $\Gamma_{M}$ of $\hat{M}$ is isomorphic to $\Gamma$. In these cases, we refer to $M$, the map $\hookrightarrow N, M^\hat{M}$, the projection $p_{\hat{M}}:\hat{M}\to M$, and $\Gamma_{M}$ as a fundamental group surjective (FS) submanifold, an FS map, an FS cover, an FS covering map, and an FS deck transformation group, respectively.
A compact core of a manifold is a compact submanifold homotopy equivalent to the ambient manifold. When a manifold is homotopy equivalent to a finite complex $K$, for $n=3$, Scott and Tucker [37] proved the existence of a compact core, while for $n-\dim K\geq 3$, Stallings [43] established this fact. Unfortunately, compact cores do not always exist (see Venema [47]). Although compact cores are preferable, we need to resort to FS submanifolds. This is fine for our purposes since we are working in coarse geometry essentially.
We will assume that $\partial M$ is convex and smooth in $N$. Let $d_{M}$ denote the path metric induced from a Riemannian metric on $N$, and let $d_{\hat{M}}$ denote the path metric on $\hat{M}$ induced from it.
  • Let $\mathbf{U}M$ denote the space of direction vectors on $M$, i.e., the space of elements of the form $(x,\vec{v})$ for $x\in M$ and a unit vector $\vec{v}\in T_{x}M-\{O\}$ with the projection $\pi_{\mathbf{U}M}:\mathbf{U}M\rightarrow M$.
  • Let $\mathbf{U}\hat{M}$ denote the space of direction vectors on $\hat{M}$ covering $\mathbf{U}M$ with the deck transformation group $\Gamma_{M}$. Let $\pi_{\mathbf{U}\hat{M}}:\mathbf{U}\hat{M}\rightarrow\hat{M}$ denote the projection.
  • Let $d_{\mathbf{U}M}$ denote the path metric on $\mathbf{U}M$ obtained from the natural Riemannian metric on the unit tangent bundle $\mathbf{U}M$ of $M$ obtained from the Riemannian metric of $M$ as a sphere bundle. We will use $d_{\mathbf{U}\hat{M}}$ to denote the induced path metric on $\mathbf{U}\hat{M}$.
A complete isometric geodesic in $\hat{M}$ is a geodesic that is an isometry of $\mathbb{R}$} into $\hat{M}$ equipped with $M$. Some but not all Riemannian geodesics are isometries. Actually, in the geodesic metric spaces, these are identical to the notion of geodesics. Since we are working on Riemannian metric spaces, we use the adjectives. A complete isometric geodesic in $M$ is a geodesic that lifts to a complete isometric geodesic in $\hat{M}$. A isometric ray in $\hat{M}$ is an isometric geodesic defined on $0,\infty)$.
[…]
Let $\rho:\pi_{1}(N)=\Gamma_{M}\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ be a homomorphism.
There is an action of $\Gamma_{M}$ on $\mathbf{UC}\hat{M}\times{\mathbb{R}}^{n}$ given by
$\gamma\cdot(x,\vec{v})=(\gamma(x),\rho(\gamma)(\vec{v}))\hbox{ for }x\in\mathbf{UC}\hat{M},\vec{v}\in{\mathbb{R}}^{n}.$
Let ${\mathbb{R}}^{n}_{\rho}$ denote the bundle over $\mathbf{UC}M$ given by taking the quotient $\mathbf{UC}\hat{M}\times{\mathbb{R}}^{n}$ by $\Gamma_{M}$.
Recall that there is a geodesic flow $\phi$ in $\mathbf{U}M$ that restricts to one in $\mathbf{UC}M$, also denoted by $\phi$. This lifts to a flow, still denoted by $\phi$, on $\mathbf{UC}\hat{M}$. Hence, there is a flow $\Phi$ on $\mathbf{UC}\hat{M}\times\mathbb{R}^{n}$, which descends to a flow, also denoted by $\Phi$, on $\mathbb{R}^{n}_{\rho}$.
1.2. Main result
In this paper, we establish connections between the partial hyperbolicity in the context of dynamical systems and the P-Anosov properties.
Unlike the abstract dominated splitting discussed in many previous articles such as [5] and [30], here we require actual splitting of smooth bundles and smooth flows, along with the presence of expanding and contracting properties in addition to the domination property.
We need the following objects since we need to work on manifolds and not just on groups.
Definition 1.1 (Partial hyperbolicity).
Suppose that $M$ is a compact Riemannian manifold with convex boundary. Let $\hat{M}$ be a regular cover of $M$ with the deck transformation group $G_{M}$. A representation $\rho:G_{M}\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ is partially hyperbolic in a bundle sense over $M$ if the following hold for a Riemannian metric on $M$:
  1. (i)
    There exist nontrivial $C^{0}$-subbundles ${\mathbb{V}}_{+},{\mathbb{V}}_{0}$, and $\mathbb{V}$$_{-}$ in ${\mathbb{R}}^{n}_{\rho}$ with fibers
    ${\mathbb{V}}_{+}(x),{\mathbb{V}}_{0}(x),{\mathbb{V}}_{-}(x)\subset{\mathbb{V}}(x)$
    for the fiber ${\mathbb{V}}(x)$ of $\mathbb{V}$} over each point $x$ of $\mathbf{UC}M$ invariant under the flow $\Phi_{t}$ of $\mathbb{V}$} over the geodesic flow $\phi_{t}$ of $\mathbf{UC}M$.
  2. (ii)
    ${\mathbb{V}}_{+}(x),{\mathbb{V}}_{0}(x)$ and ${\mathbb{V}}_{-}(x)$ for each $x\in\mathbf{UC}M$ are independent subspaces, and their sum equals ${\mathbb{V}}(x)$.
  3. (iii)
    For any fiberwise metric on ${\mathbb{R}}^{n}_{\rho}$ over $\mathbf{UC}M$, the lifted action of $\Phi_{t}$ on $\mathbb{V}$$_{+}$ (resp. $\mathbb{V}$$_{-}$) is dilating (resp. contracting). That is, for coefficients $A>0,a>0$, $A^{\prime}>0,a^{\prime}>0$:
    1. (a)
[…]
    2. (b)
[…]
    3. (c)
      (Dominance properties)
[…]
We additionally require that $\dim{\mathbb{V}}_{+}=\dim{\mathbb{V}}_{-}\geq 1$. Here $\dim{\mathbb{V}}_{+}$ is the partial hyperbolicity index of $\rho$ for the associated bundle decomposition ${\mathbb{V}}_{+},{\mathbb{V}}_{0},{\mathbb{V}}_{-}$ where $\dim{\mathbb{V}}_{+}$ is maximal and $\dim{\mathbb{V}}_{+}=\dim{\mathbb{V}}_{-}<n/2$, so that $\dim{\mathbb{V}}_{0}\geq 1$. Furthermore, ${\mathbb{V}}_{0}$ is said to be the neutral subbundle of $\mathbb{V}$}. (See Definition 1.5 of [10] or [6]. We use the words “in a bundle sense” only as a convenience of this paper.)
The definition depends on the Riemannian metric, but not on $\left|\mkern-2.0mu\left|\cdot\right|\mkern-2.0mu\right|_{{\mathbb{R}}^{n}_{\rho}}$. However, if $\Gamma_{M}$ is word hyperbolic, then we can replace $\mathbf{UC}\hat{M}$ with a Gromov flow space which depends only on $\Gamma_{M}$ up to quasi-isometry. (See Theorem 2.6 and Gromov [23] or Mineyev [36].)
We say that $N$ has a partially hyperbolic linear holonomy group in a bundle sense if
  • there exists an FS submanifold $M$ in $N$ with convex boundary for a Riemannian metric on $N$, and
  • the linear part $\rho:\pi_{1}(N)=\Gamma_{M}\to{\mathsf{GL}}(n,\mathbb{R})$ is a partially hyperbolic representation in a bundle sense over $M$.
Since partial hyperbolicity is stronger than the Anosov property by Theorem 1.1, $\pi_{1}(N)$ is always hyperbolic. By Theorem 2.6, there exists a quasi-isometry from $\mathbf{UC}\hat{M}$ to the Gromov flow space $\partial^{(2)}_{\infty}M\times\mathbb{R}$ preserving the geodesic length parameter. Hence, the discussion reduces to one on the Gromov flow space, which is independent of the choice of FS submanifolds.
We denote by $a_{1}(A),\ldots,a_{n}(A)$ the nonincreading set of singular values of an $n\times n$-matrix $A$. We denote by $\lambda_{1}(A),\ldots,\lambda_{n}(A)$ the nonincreasing set of norms of eigenvalues of an $n\times n$-matrix $A$. (Here, the multiplicity is allowed.)
Let $\log\lambda_{1},\dots,\log\lambda_{n}$ denote the logarithmic eigenvalue functions defined on the maximal Abelian Lie algebra of $\mathfrak{gl}(n,{\mathbb{R}})$. For roots
$\theta=\left\{\log\lambda_{i_{1}}-\log\lambda_{i_{1}+1},\ldots,\log\lambda_{i_{m}}-\log\lambda_{i_{m}+1}\right\},1\leq i_{1}<\cdots<i_{m}\leq n-1,$
of $\mathrm{GL}(n,{\mathbb{R}})$, the parabolic subgroup $P_{\theta}$ (resp. $\left.P_{\theta}^{-}\right)$ is the subgroup of block upper (resp. lower) triangular matrices in $\mathrm{GL}(n,{\mathbb{R}})$ with square diagonal blocks of sizes $i_{1},i_{2}-i_{1},\ldots,i_{m}-i_{m-1},n-i_{m}$ respectively.
Theorem 1.1.
Let $N$ be an $m$-manifold with a finitely generated fundamental group. Let $\rho:\pi_{1}(N)\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ be a representation. Assume one of the following:
  • $|\log\lambda_{i_{g}}(\rho(g))|=1$ for all $g\in\pi_{1}(M)$ where the index $i_{g}\in\{1,\dots,n\}$ may depend on $g$.
  • There exists a constant $C>0$ such that $|\log a_{j_{g}}(\rho(g))|<C$ for all $g\in\pi_{1}(M)$ where the index $j_{g}\in\{1,\dots,n\}$ may depend on $g$.
Then $\rho$ is $P$-Anosov for any parabolic subgroup $P$ of ${\mathsf{GL}}(n,{\mathbb{R}})$ with $P=P_{\theta}$ for $\theta$ containing $\log\lambda_{k}-\log\lambda_{k+1},k\leq n/2$ if and only if $\rho$ is partially hyperbolic of the index $k$ in a bundle sense. In these cases $k<n/2$.
The significance of this result is that we work with the full general linear group ${\mathsf{GL}}(n,\mathbb{R})$ instead of orthogonal groups. Also, Anosov representations correspond to only dominated representations and not actual partially hyperbolic flows as here. Promoting the dominated representation to Anosov ones is trivial for some Lie groups but not for general linear groups.
We can ask if an Anosov representation is always partially hyperbolic up to multiplying by a positive scalar representation to $\mathbb{R}$}. We cannot answer this question since controlling the singular values of representation elements is very hard in general, and the only related results that we know of are due to Abel-Margulis-Soifer [1], which is not applicable here as far as we know.
Corollary 1.2.
Suppose that $N$ is a complete affine $n$-manifold with a finitely generated fundamental group. Let $\rho:\pi_{1}(N)\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ be a linear part of the affine holonomy group. Then the conclusions of Theorem 1.1 hold.
In the field of affine manifolds, the hyperbolic flow was already discovered by Goldman and Labourie [21] for $3$-dimensional Margulis space-times, and Ghosh [15], Ghosh-Treib [19], and Danciger-Zhang [13] prove these properties for affine actions on higher-dimensional affine spaces. However, the linear parts are always assumed to be in the orthogonal groups or Hitchin, where the properties are more or less immediate.
We wish to approach the conjecture that a closed complete affine manifold cannot have a word hyperbolic fundamental group. One approach considered is using Osedelec-type theorems (See Chapter 8 of [31]). However, generalizing our results to measurable settings is currently hindered by numerous technical issues. Discussions on these matters took place with researchers such as Nicholas Tholozan at BIRS, Banff, during the winter of 2020.
1.3. Some historical comments on complete affine manifolds
Originally, Milnor posed the question whether a free affine group of rank $\geq 2$ could act freely and properly discontinuously on affine space. Margulis provided an affirmative answer by constructing examples in the affine $3$-space, where the linear parts belong to ${\mathsf{SO}}(2,1)$. Drumm and Goldman subsequently explained this result using fundamental domains in a series of papers (see [11] for a survey). Sullivan [44] observed that the linear part of each affine group element must have $1$ as an eigenvalue and proved the Chern conjecture for complete affine manifolds [33]. Klingler [32] later extended this to closed special affine manifolds.
Further developments include Ghosh [15, 17, 16, 18], Ghosh-Treib [19], and Smilga [38, 39, 40, 41, 42], who generalized these results and produced higher-dimensional examples of proper and free affine actions. Abels, Margulis, and Soifer also contributed to related topics such as the Auslander conjecture (see Abels [2] for a survey). Danciger and Zhang [13] investigated affine deformations of Hitchin representations, showing that they do not exist. Danciger, Kassel, and Guéritaud [12] provided first examples of fundamental groups of closed manifolds acting properly and freely on affine spaces.
Although these works are related to the themes of this paper, their specific aims differ. This summary only scratches the surface of this rapidly evolving field.
1.4. Outline
In Section 2, we will introduce affine structures and some basic facts on $\delta$-hyperbolic metric spaces. Suppose that $\Gamma_{M}$ is word-hyperbolic. Then $\hat{M}$ has a Gromov hyperbolic Riemannian metric by the Svarc-Milnor theorem. (see [35] and [45]. Svarc was a translated name of Albert Schwarz who is currently working as a professor in the University of California, Davis. He published this theorem earlier than Milnor in Russian.) We discuss the ideal boundary of $\hat{M}$, identifiable with the Gromov boundary, and the complete isometric geodesics in $\hat{M}$. We relate the space of complete isometric geodesics to the Gromov flow space.
In Section 3, we provide a review of the definition of the $P$-Anosov property of Guichard-Wienhard [25]. For convenience for reductive groups, we will use the approaches of Gúeritaud-Guichard-Kassel-Wienhard [24] and Bochi-Potrie-Sambarino [5] using singular values of the linear holonomy group. (The theory of Kapovich, Leeb, and Porti [27], [29] is equivalent to these, but it is not so tailored for our purposes here.) We give various associated definitions.
Section 4 has the main arguments. In Section 4.1, we will relate the partial hyperbolic property in the singular-value sense to that in the bundle sense. In Section 4.2, we show that the linear part of the holonomy representation of a complete affine manifold is $P$-Anosov if and only if it is partially hyperbolic. First, we do this when the linear holonomy group has a reductive Zariski closure. This involves the fact that $1$ has to be an eigenvalue of each element of the linear holonomy groups of complete affine manifolds. We relate it to singular values using the clever ideas of Danciger and Stecker, the spectrum theory of Breulliard-Sert [7], Benoist [3], [4], and the work of Potrie-Kassel [30]. We could generalize to the cases of nonreductive linear holonomy groups using the stability of the Anosov and partial hyperbolic properties. This will prove Theorem 1.1.
1.5. Acknowledgments
We thank Michael Kapovich for various help with geometric group theory and coarse geometry. This article began with some discussions with Michael Kapovich during the conference honoring the 60th birthday of William Goldman at the University of Maryland, College Park, in 2016. We thank Jeffrey Danciger and Florian Stecker for help with the $P$-Anosov properties, which yielded the main idea for the proof of Theorem 1.1. We also thank Herbert Abels, Richard Canary, Virginie Charette, Todd Drumm, William Goldman, François Guéritaud, Fanny Kassel, Andrés Sambarino, Nicholas Tholozan, and Konstantinos Tsouvalas for various discussions helpful for this paper. We also thank BIRS, Banff, Canada, and KIAS, Seoul, where some of this work was done. This paper is a joint work with Kapovich in the beginning and Danciger and Stecker later. Unfortunately, they did not wish to join as coauthors for various reasons.
2. Preliminary
We recall some standard well-known facts in this section to set the notations and so on. Also, we need a slight modification of the theories for our purposes. Most of the background on geometric groups and coarse geometry is from the comprehensive textbook by Druţu and Kapovich [14].
The other purpose is to establish the Anosov properties for manifolds using coarse isometries since most of the literature works on with groups.
2.1. Convergences of geodesics
Let $M$ be a manifold of dimension $n$ with a regular cover $\hat{M}$ and the deck transformation group $\Gamma_{M}$. We assume that $M$ has a Riemannian metric with convex boundary.
We say that a sequence of isometric rays (resp. complete isometric geodesics) $l_{i}$ converges to an isometric ray (resp. complete geodesic) $l$ (denoted as $l_{i}\to l$) if $l_{i}(t)\to l(t)$ for each $t\in[0,\infty)$ (resp. $t\in\mathbb{R}$).
By the Arzelà-Ascoli theorem and the properness of $\hat{M}$, any sequence of isometric rays (resp. complete isometric geodesics) converges to an isometric ray (resp. complete isometric geodesic) if the sequence of pairs of their 0-points and directions at the 0-points converges in $\mathbf{U}\hat{M}$. (See Section 11.11 of [14].)
2.2. Gromov hyperbolic spaces
Let us recall some definitions: A metric space is geodesic if every pair of points are connected by a geodesic, i.e., a path that is an isometry from an interval to the metric space. We will only work with geodesic metric spaces. Let $(X,d_{X})$ and $(Y,d_{Y})$ be metric spaces. A map $f:X\rightarrow Y$ is called $(L,C^{\prime})$-coarse Lipschitz for $L,C>0$ if
$d_{Y}(f(x),f(x^{\prime})\leq Ld_{X}(x,x^{\prime})+C^{\prime}\hbox{ for all }x,x^{\prime}\in X.$
A map $f:X\rightarrow Y$ is an $(L,C)$-quasi-isomeric embedding if $f$ satisfies
$L^{-1}d_{X}(x,x^{\prime})-C\leq d_{Y}(f(x),f(x^{\prime}))\leq Ld_{X}(x,x^{\prime})+C\hbox{ for all }x,x^{\prime}\in X.$
A map $\bar{f}:Y\rightarrow X$ is a $C$-coarse inverse of $f$ for $C>0$ if
$d_{X}(\bar{f}\circ f,{\rm I}_{X})\leq C,\hbox{ and }d_{Y}(f\circ\bar{f},{\rm I}_{Y})\leq C.$
A map $f:X\rightarrow Y$ between metric spaces is called a quasi-isometry if it is a coarse Lipschitz map and admits a coarse Lipschitz coarse inverse map. (Note that these are not necessarily continuous, as we follow Druţu-Kapovich [14].)
Lemma 2.1.
Let $M$ be a compact Riemannian manifold with possibly nonempty boundary. $\pi_{\mathbf{U}\hat{M}}$ is a quasi-isometry.
Proof.
Each fiber is uniformly bounded under $d_{\mathbf{U}\hat{M}}$. ∎
A metric space $(X,d)$ is proper if every metric ball has a compact closure. In other words, $d_{p}(\cdot)=d(p,\cdot)$ is a proper function $X\rightarrow{\mathbb{R}}$. Clearly, a complete Riemannian manifold is a proper metric space. We denote by $B^{d}(x,R)$ the ball of radius <R<R in $X$ with the center $x\in X$.
Let $X$ be a geodesic metric space with metric $d$. A geodesic triangle $T$ is a concatenation of three geodesics $\tau_{1},\tau_{2},\tau_{3}$ where the indices are modulo $3$. The thinness radius of a geodesic triangle $T$ is the number
$\delta(T):=\max_{j=1,2,3}\left(\sup_{p\in\tau_{j}}d(p,\tau_{j+1}\cup\tau_{j+2})\right).$
A geodesic metric space $X$ is called $\delta$-hyperbolic in the sense of Rips if every geodesic triangle $T$ in $X$ is $\delta$-thin.
By Corollary 11.29 of [14] as proved by Section 6.3C of Gromov [23], the Gromov hyperbolicity is equivalent to the Rips hyperbolicity for geodesic metric spaces. We will use these concepts interchangeably.
We will assume that $\Gamma_{M}$ is word-hyperbolic for our discussions below. This is equivalent to assuming that $\hat{M}$ is Gromov hyperbolic by the Svarc-Milnor lemma.
Two isometric rays $\rho_{1}$ and $\rho_{2}$ of $X$ are equivalent if $t\mapsto d(\rho_{1}(t),\rho_{2}(t))$ is bounded. This condition is equivalent to the one that their Hausdorff distance under $d$ is bounded. Assume that $X$ is $\delta$-hyperbolic. $X$ has a well-defined ideal boundary $\partial_{\infty}X$ as in Definition 3.78 of [14]; that is, the space of the equivalence classes of isometric rays. (See [9] and [20].) We denote by $\partial^{(2)}_{\infty}X=\partial_{\infty}X\times\partial_{\infty}X-\Delta$ where $\Delta$ is the diagonal set.
The constant $C$ below is called a quasi-geodesic constant.
Lemma 2.2.
Let $M$ be a compact manifold with the induced path metric $d_{\hat{M}}$ on $\hat{M}$ from a Riemannian metric of $M$. Let $\Gamma_{M}$ be the deck transformation group of $\hat{M}\rightarrow M$. Suppose that $\Gamma_{M}$ is word-hyperbolic.
Then there exists a constant $C>0$ so that for every pair of complete isometric geodesics $l_{1}$ and $l_{2}$ in $\hat{M}$ with the identical set of endpoints in $\partial_{\infty}\hat{M}$, the Hausdorff distance between $l_{1}$ and $l_{2}$ under $d$ is bounded above by $C$.
Proof.
Since $l_{1}$ and $l_{2}$ are both $(1,0)$-quasigeodesics, this follows from Proposition 3.1 of Chapter 3 of [9]. ∎
We put on $\hat{M}\cup\partial_{\infty}\hat{M}$ the shadow topology, which is a first-countable Hausdorff topology by Lemma 11.76 of [14]. It is also compact according to Section 11.11 of [14]. Since it is first countable, we do not need to consider nets on the space but only sequences to understand the continuity of the real-valued functions.
We denote by $\partial_{G}X$ the Gromov boundary of a $\delta$-hyperbolic geodesic metric space $X$. (See Section 11.12 of [14].) By Theorem 11.104 of [14], there is a homeomorphism $h:\partial_{\infty}X\rightarrow\partial_{G}X$ given by sending the equivalence class $\rho]$ of an isometric ray $\rho$ to the equivalence class of $\{\rho(n)\}_{n\in{\mathbb{Z}}_{+}}$. We will identify these two spaces using this map.
Let $\hat{\Gamma}_{M}$ denote the Cayley graph of $\Gamma_{M}$. $\hat{\Gamma}_{M}$ and $\hat{M}$ are quasi-isometric by the Svarc-Milnor lemma again. The Gromov boundary $\partial_{\infty}\hat{M}$ of $\hat{M}$ and can be identified with the boundary $\partial_{\infty}\Gamma$ of $\hat{\Gamma}_{M}$ with the word metric by Theorem 11.108 of [14].
For any Gromov hyperbolic geodesic metric space $X$, the set of $GX$ of complete isometric geodesics has a metric: for $g,h\in GX$, we define the metric $d_{GX}$ given by
(2) $d_{GX}(g,h):=\int^{\infty}_{-\infty}d_{X}(g(t),h(t))2^{-|t|}dt.$
Notice that $d_{X}(g(t),h(t))\leq 2|t|$ since we can follow the geodesic $g$ backward and then forward in $h$ to obtain a distance $2|t|$. Thus, the above integral is always well defined. (See Gromov [23].) Let $GX$ have this metric topology. (See Section 11.11 of [14].)
Lemma 2.3.
Suppose that $\Gamma_{M}$ is Gromov hyperbolic. Then there is a quasi-isometric homeomorphism $\mathcal{F}:\mathbf{UC}\hat{M}\rightarrow G\hat{M}$ by taking a unit vector $\vec{u}$ at $x$ to a complete isometric geodesic $^{\mathbb{R}}\rightarrow\hat{M}$ passing $x$ tangent to $\vec{u}$.
Proof.
There is a map $G\hat{M}\to\mathbf{UC}\hat{M}$ given by sending the complete isometric geodesic $g:\mathbb{R}\to\hat{M}$ to $(g(0),\vec{v}_{0})$ where $\vec{v}_{0}$ is the unit tangent vector at $g(0)$. This map is clearly bijective. In Section III of [34], the map $G\hat{M}\to\hat{M}$ defined by $g\mapsto g(0)$ is shown to be a quasi-isometry. Therefore, post-composing this map with a lift that sends each $g(0)$ to a vector in the fiber of $\mathbf{UC}\hat{M}$ over $g(0)$ results in a quasi-isometry. For a sequence $\{g_{i}\}$ of geodesics in $\hat{M}$, if $d_{GX}(g,g_{i})\to 0$, then $(g_{i}(0),g_{i}^{\prime}(0))\to(g(0),g^{\prime}(0))$ since otherwise, we would obtain a positive lower bound for the integral (2).
The inverse map $\mathbf{UC}\hat{M}\to G\hat{M}$ is also continuous due to the continuity of the exponential map and considerations involving (2), where $d_{X}(g(t),h(t))$ grows sublinearly. The continuity of the integral values under $g$ and $h$ can be shown by cutting off for $|t|>N$ where $N$ is independent of $g$ or $h$ by (2) and an estimate.
2.3. Flow space $G\hat{M}$ quasi-isometric to $\partial_{\infty}^{(2)}\hat{M}\times{\mathbb{R}}$
We assume that $\Gamma_{M}$ is word-hyperbolic. For a word-hyperbolic group $\Gamma$, we let $\hat{\Gamma}$ denote its Cayley graph. For simplicity, we denote by $\partial_{\infty}\Gamma$ the boundary $\partial_{\infty}\hat{\Gamma}$. We denote
$\partial_{\infty}^{(2)}\Gamma:=\partial_{\infty}\Gamma\times\partial_{\infty}\Gamma-\{(t,t)|t\in\partial_{\infty}\Gamma\}.$
We denote by $F\Gamma_{M}$ the Gromov geodesic flow space of $\hat{\Gamma}_{M}$ obtained by the following proposition. (Note that we identified $\partial_{\infty}\hat{M}$ with $\partial_{\infty}\Gamma$ in Section 2.2.)
Theorem 2.4 (Theorem 8.3.C of Gromov [22], Theorem 60 of Mineyev [36]).
Let $\Gamma$ be a finitely generated word hyperbolic group.
Then there exists a proper hyperbolic metric space $F\Gamma$ with the following properties:
  • (i)
    $\Gamma\times{\mathbb{R}}\times{\mathbb{Z}}/2{\mathbb{Z}}$ acts on $F\Gamma$.
  • (ii)
    The $\Gamma\times{\mathbb{Z}}/2{\mathbb{Z}}$-action is isometric.
  • (iii)
    Every orbit $\Gamma\rightarrow F\Gamma$ is a quasi-isometry.
  • (iv)
    The $\mathbb{R}$}-action is free, and every orbit ${\mathbb{R}}\rightarrow F\Gamma$ is a quasi-isometric embedding. The induced map $F\Gamma/{\mathbb{R}}\rightarrow\partial_{\infty}^{(2)}\Gamma$ is a homeomorphism.
We shall say that $F\Gamma$ is the flow space of $\Gamma$. In fact, $F\Gamma$ is unique up to a $\Gamma\times{\mathbb{Z}}/2{\mathbb{Z}}$-equivariant quasi-isometry sending $\mathbb{R}$}-orbits to $\mathbb{R}$}-orbits. We shall denote by $\phi_{t}$ the $\mathbb{R}$}-action on $F\Gamma$ and by $(\tau_{+},\tau_{-}):F\Gamma\rightarrow F\Gamma/{\mathbb{R}}\cong\partial_{\infty}^{(2)}\Gamma$ the maps associating to an element of $F\Gamma$ the endpoint of its $\mathbb{R}$}-orbit. Gromov identifies $F\Gamma$ with $\partial_{\infty}^{(2)}\Gamma\times{\mathbb{R}}$ with a certain metric called the Gromov metric.
Proposition 2.5.
Let $M$ be a compact manifold with a covering map $\hat{M}\rightarrow M$ with a deck transformation group $\Gamma_{M}$. Suppose that $\Gamma_{M}$ is word-hyperbolic.
Then there is a quasi-isometric surjective map $\mathcal{E}:G\hat{M}\rightarrow\partial^{(2)}_{\infty}\Gamma_{M}\times{\mathbb{R}}$ with compact fibers. The map obtained from composing with a projection to the first factor is a map given by taking the endpoints of complete isometric geodesics.
Proof.
Choose arbitrarily a base point $x_{0}$ in $\hat{M}$. For each complete isometric geodesic $g:{\mathbb{R}}\rightarrow\hat{M}$, let $g_{x_{0}}$ denote a projection point on $g({\mathbb{R}})$ that is of the minimal distance from $x_{0}$ and let $\partial_{+}g,\partial_{-}g\in\partial_{\infty}\Gamma_{M}$ denote the forward and backward endpoints, and let $t(g)\in{\mathbb{R}}$ denote $\pm d(g(0),g_{x_{0}})$ where we use ++ if $g(0)$ is ahead of $g_{x_{0}}$.
We define $\mathcal{E}(g)=(\partial_{+}g,\partial_{-}g,t(g))$. The surjectivity follows from Proposition 2.1 of Chapter 2 of [9] since $t$ is an isometry on each complete isometric geodesic. The compactness of the fiber is implied by Lemma 2.2.
Champetier [8] constructs the space $G\hat{\Gamma}_{M}$ from $\hat{\Gamma}_{M}$ quasi-isometric to $\Gamma_{M}$ following Gromov. Recall from Section 2.2 that $\partial_{\infty}\hat{M}=\partial_{\infty}\Gamma_{M}$. Using the orbit map $\Gamma_{M}\to\hat{M}$ sending $\gamma$ to $\gamma(x_{0})$, we extend the map to $\hat{\Gamma}_{M}\to\hat{M}$ by sending each edge to a shortest geodesic.
Let $I:G\hat{\Gamma}_{M}\to G\hat{M}$ be a map that sends a geodesic $g$ in $\hat{\Gamma}_{M}$ to a complete isometric geodesic $g^{\prime}$ in $\hat{M}$ with the same endpoints in $\partial_{\infty}\hat{M}$, and maps $g(0)$ to the nearest point on the image of $g^{\prime}$.
The map $I^{\prime}:G\hat{M}\to G\hat{\Gamma}_{M}$ can be defined by taking a complete isometric geodesic $g$ in $\hat{M}$ to a geodesic $g^{\prime}$ in $\hat{\Gamma}_{M}$ in a similar manner, with $g(0)$ going to one of the elements of $\Gamma_{M}(x_{0})$ nearest to it.
By Theorem 11.72 of [14], which states that every isometric geodesic in $\hat{\Gamma}_{M}$ is a quasi-geodesic in $\hat{M}$ and vice versa, and considering that every complete isometric geodesic in $\hat{M}$ is uniformly bounded away from one in $\hat{\Gamma}_{M}$ in the Hausdorff distance, it follows that $G\hat{M}$ is quasi-isometric to $G\hat{\Gamma}_{M}$ via the maps $I$ and $I^{\prime}$, using (2).
We define $\mathcal{E}^{\prime}:G\hat{\Gamma}_{M}\rightarrow\partial_{\infty}^{(2)}\Gamma_{M}\times{\mathbb{R}}$ as we did for $\mathcal{E}$ above. By Proposition 4.8 and (4.3) of [8], $\mathcal{E}^{\prime}$ sends $G\hat{\Gamma}_{M}$ to $\partial_{\infty}^{(2)}\Gamma_{M}\times{\mathbb{R}}=\partial_{\infty}^{(2)}\hat{M}\times{\mathbb{R}}$ as a quasi-isometric homeomorphism. Since $\mathcal{E}(g,t)$ and $\mathcal{E}^{\prime}\circ I^{\prime}(g,t)$ are uniformly bounded away from each other, the result follows. ∎
The following shows that $\mathbf{UC}\hat{M}$ can be used as our Gromov flow space.
Theorem 2.6.
The map $\mathcal{E}\circ\mathcal{F}:\mathbf{UC}\hat{M}\rightarrow\partial_{\infty}^{(2)}\Gamma_{M}\times{\mathbb{R}}$ is a quasi-isometry where each complete isometric geodesic in $\mathbf{UC}\hat{M}$ goes to $(t_{1},t_{2})\times{\mathbb{R}}$ for its pair $(t_{1},t_{2})\in\partial_{\infty}^{(2)}\Gamma_{M}$ of endpoints.
Proof.
Proposition 2.5 and Lemma 2.3 imply the result. ∎
3. Partial hyperbolicity and $P$-Anosov properties
The purpose of this section is to set notations and recall standard facts since some of these are not too well-known to many geometric topologists not in this specific area.
3.1. Some background
We have to repeat some of the standard materials from Section 2.2 of [24] and Chapter 2 of [26]. Consider a real reductive Lie group $G$ with a maximal compact subgroup $K$.
We assume that $\mathrm{Ad}(G)\subset{\mathbf{Aut}}(\mathfrak{g})$ for simplicity. For example, this holds if $G$ is connected. We can usually reduce everything here to the case where $G$ is connected. Here, $G$ is an almost product of $Z(G)_{0}$ and $G_{s}$ where $Z(G)_{0}$ is the identity component of the center $Z(G)$ of $G$ and $G_{s}=D(G)$ the derived subgroup of $G$, which is semisimple.
A parabolic subgroup $P$ of $G$ is a subgroup $P$ of the form $G\cap\mathbf{P}({\mathbb{R}})$ for some algebraic subgroup $\mathbf{P}$ of an algebraic group $\mathbf{G}$ where $\mathbf{G}({\mathbb{R}})/\mathbf{P}({\mathbb{R}})$ is compact and $\mathbf{G}({\mathbb{R}})=G$.
Let $\mathfrak{z}(\mathfrak{g})$ denote the Lie algebra of $Z(G)$, and $\mathfrak{g}_{s}$ the Lie algebra of $G_{s}$. Then $\mathfrak{g}=\mathfrak{z}(\mathfrak{g})\oplus\mathfrak{g}_{s}$. Letting $\mathfrak{t}$ denote the Lie algebra of $K$, we have $\mathfrak{g}=\mathfrak{t}\oplus\mathfrak{q}$ where $\mathfrak{q}$ is the 1-1-eigenspace of an involution $\theta$ preserving $\mathfrak{t}$ as the $1$-eigenspace. Let $\mathfrak{a}\subset\mathfrak{q}$ be the maximal abelian subalgebra, i.e., the Cartan subalgebra of $\mathfrak{g}$. Then $\mathfrak{a}=\mathfrak{z}(\mathfrak{g})\cap\mathfrak{q}\oplus\mathfrak{a}_{s}$ where $\mathfrak{a}_{s}=\mathfrak{a}\cap\mathfrak{g}_{s}$, a maximal Abelian subalgebra of $\mathfrak{q}\cap\mathfrak{g}_{s}$.
We have a decomposition to $ad(\mathfrak{a})$-eigenspaces $\mathfrak{g}=\mathfrak{g}_{0}\oplus\bigoplus_{\alpha\in\Sigma}\mathfrak{g}_{\alpha}$. Here, $\mathfrak{g}_{0}$ is the direct sum of $z(\mathfrak{g})$ and the centralizer of $\mathfrak{a}$ in $\mathfrak{g}_{s}$. The set $\Sigma\subset\mathfrak{a}^{*}=\mathrm{Hom}_{\mathbb{R}}(\mathfrak{a},\mathbb{R})$ projects to a (possibly nonreduced) root system of $\mathfrak{a}_{s}^{*},$ and each $\alpha\in\Sigma$ is called a restricted root of $\mathfrak{a}$ in $\mathfrak{g}$. There is a subset $\Delta\subset\Sigma$ called a simple system, i.e. a subset such that any root is expressed uniquely as a linear combination of elements of $\Delta$ with coefficients all of the same sign; the elements of $\Delta$ are called simple roots. There is a subset $\Sigma^{+}\subset\Sigma$ which is the set of positive roots, i.e. roots that are nonnegative linear combinations of elements of $\Delta$, where $\Sigma=\Sigma^{+}\cup\left(-\Sigma^{+}\right)$ holds additionally.
We define $P_{\theta}$ as the parabolic subgroup with the Lie algebra
$\operatorname{Lie}\left(P_{\theta}\right)=\mathfrak{g}_{0}\oplus\bigoplus_{\alpha\in\Sigma^{+}}\mathfrak{g}_{\alpha}\oplus\bigoplus_{\alpha\in\Sigma^{+}\backslash\Sigma_{\theta}^{+}}\mathfrak{g}_{-\alpha}.$
The connected components of $\mathfrak{a}\setminus\cup_{\alpha\in\Sigma}\ker(\alpha)$ are called Weyl chambers of $\mathfrak{a}$. The component where every $\alpha\in\Sigma_{+}$ is positive is denoted by $\mathfrak{a}^{+}$.
Let $G={\mathsf{SL}}_{\pm}(n,{\mathbb{R}})$. The maximal abelian subalgebra $A_{n}$ of $\mathfrak{p}$ in this case is the subspace of diagonal matrices with entries $a_{1},\dots,a_{n}$ where $a_{1}+\dots+a_{n}=0$. Let $\log\lambda_{i}:A_{n}\rightarrow{\mathbb{R}}$ denote the projection to the $i$-th factor. $\Sigma$ consists of $\alpha_{ij}:=\log\lambda_{i}-\log\lambda_{j}$ for $i\neq j$, and $\mathfrak{sl}(n,{\mathbb{R}})_{\alpha_{ij}}=\{cE_{ij}|c\in{\mathbb{R}}\}$ where $E_{ij}$ is a matrix with entry $1$ at $(i,j)$ and zero entries everywhere else.
A positive Weyl chamber is given by
$A_{n}^{+}=\{(a_{1},\dots,a_{n})|a_{1}\geq\dots\geq a_{n},a_{1}+\cdots+a_{n}=0\},$
and we are given
[…]
$\Delta=\{\alpha_{i}:=\log\lambda_{i}-\log\lambda_{i+1}|i=1,\dots,n-1\}.$
We recall Example 2.14 of [24]: Let $G$ be $\mathrm{GL}(n,{\mathbb{R}})$ seen as a real Lie group. Its derived group is $G_{s}=D(G)=\mathrm{SL}(n,{\mathbb{R}}).$ We can take $\mathfrak{a}\subset\mathfrak{gl}(n,{\mathbb{R}})$ to be the set of real diagonal matrices of size $n\times n$. As above, for $1\leq i\leq n,$ let $\log\lambda_{i}\in\mathfrak{a}^{*}$ be the evaluation of the $i$-th diagonal entry. Then
$\mathfrak{a}=\mathfrak{z}(\mathfrak{g})\cap\mathfrak{a}\oplus\mathfrak{a}_{s}\hbox{ where }\mathfrak{z}(\mathfrak{g})\cap\mathfrak{a}=\bigcap_{1\leq i,j\leq n}\operatorname{Ker}\left(\log\lambda_{i}-\log\lambda_{j}\right)$
is the set of real scalar matrices, and $\mathfrak{a}_{s}=\operatorname{Ker}\left(\log\lambda_{1}+\cdots+\log\lambda_{n}\right)$ is the set of traceless real diagonal matrices. The set of restricted roots of $\mathfrak{a}$ in $G$ is
$\Sigma=\left\{\log\lambda_{i}-\log\lambda_{j}\mid 1\leq i\neq j\leq n\right\}.$
We can take $\Delta=\left\{\log\lambda_{i}-\log\lambda_{i+1}\mid 1\leq i\leq n-1\right\}$ so that
$\Sigma^{+}=\left\{\log\lambda_{i}-\log\lambda_{j}\mid 1\leq i<j\leq n\right\},$
and $\mathfrak{a}^{+}$ is the set of elements of $\mathfrak{a}$ whose entries are increasing. Let ${\rm Cl}({\mathfrak{a}}^{+})$ denote its closure. For
(3) $\theta=\left\{\log\lambda_{i_{1}}-\log\lambda_{i_{1}+1},\ldots,\log\lambda_{i_{m}}-\log\lambda_{i_{m}+1}\right\}\\ \hbox{ with }1\leq i_{1}<\cdots<i_{m}\leq n-1,$
the parabolic subgroup $P_{\theta}$ (resp. $\left.P_{\theta}^{-}\right)$ is the group of block upper ( resp. lower) triangular matrices in $\mathrm{GL}(n,{\mathbb{R}})$ with square diagonal blocks of sizes $i_{1},i_{2}-i_{1},\ldots,i_{m}-i_{m-1},n-i_{m}$. In particular, $P_{\Delta}$ is the group of upper triangular matrices in $\mathrm{GL}(n,({\mathbb{R}})$.
From now on, we will fix $\mathfrak{a}_{n}\subset{\mathsf{GL}}(n,{\mathbb{R}})$ as the set of real diagonal matrices of size $n\times n$. We will also regard $\theta$ as a subset of $\{1,\dots,n-1\}$ where each $i$ corresponds to $\log\lambda_{i}-\log\lambda_{i+1}$. Also, let $\mathfrak{a}_{n}^{+}$ denote the open positive Weyl chamber.
3.2. $P$-Anosov representations
Let $(P^{+},P^{-})$ be a pair of opposite parabolic subgroups of ${\mathsf{GL}}(n,{\mathbb{R}})$, and let $\mathcal{F}^{{\pm}}$ denote the flag spaces ${\mathsf{GL}}(n,{\mathbb{R}})/P_{\pm}$. Let $\chi$ denote the unique open ${\mathsf{GL}}(n,{\mathbb{R}})$-orbit of the product $\mathcal{F}^{+}\times\mathcal{F}^{-}$. The product subspace $\chi$ has two ${\mathsf{GL}}(n,{\mathbb{R}})$-invariant distributions $E^{\pm}:=T_{x_{\pm}}\mathcal{F}^{\pm}$ for $(x^{+},x^{-})\in\chi$. Let $\phi_{t}$ be given as in Theorem 2.4 as the $\mathbb{R}$}-action on $F\Gamma_{M}$ whose orbits are quasi-geodesics. We denote by $(\tau_{+},\tau_{-}):F\Gamma_{M}\rightarrow F\Gamma_{M}/{\mathbb{R}}\cong\delta_{\infty}\Gamma_{M}^{(2)}$ the maps associating to a point the end points of its $\mathbb{R}$}-orbit.
Definition 3.1 (Definition 2.10 of [25]).
A representation $\rho:\Gamma_{M}\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ is $(P^{+},P^{-})$-Anosov if there exist continuous $\rho$-equivariant maps $\xi^{+}:\partial_{\infty}\Gamma_{M}\rightarrow{\mathcal{F}}^{+}$, $\xi^{-}:\partial_{\infty}\Gamma_{M}\rightarrow{\mathcal{F}}^{-}$ such that:
  1. (i)
    for all $(x,y)\in\partial_{\infty}^{(2)}\Gamma_{M}$, the pair $(\xi^{+}(x),\xi^{-}(y))$ is transverse, and
  2. (ii)
    for one (and hence any) continuous and equivariant family of norms $(\left|\mkern-2.0mu\left|\cdot\right|\mkern-2.0mu\right|_{m})_{m\in F\Gamma_{M}}$ on
    $(T_{\xi^{+}(\tau^{+}(m))}{\mathcal{F}}^{+})_{m\in F\Gamma_{M}}\Big(\hbox{resp. }(T_{\xi^{-}(\tau^{-}(m))}{\mathcal{F}}^{-})_{m\in F\Gamma_{M}}\Big):$
[…]
We call $\xi^{\pm}:\partial_{\infty}F\Gamma_{M}\rightarrow{\mathcal{F}}^{\pm}$ the Anosov maps associated with $\rho:\Gamma_{M}\rightarrow G$. (See also Section 2.5 of [24] for more details.)
It is sufficient to consider only the case where $P^{+}$ is conjugate to $P^{-}$ by Lemma 3.18 of [25]. (See also Definition 5.62 of [28]). In this case, $\rho$ is called $P$-Anosov for $P=P^{+}$. Note that this means that
(4) $\theta=\theta^{*}$
where $\theta^{*}$ is the image $\theta$ under the map $i\mapsto n-i$, $i=1,\dots,n-1$. Hence, this will always be true for the set $\theta$ occurring for $P_{\theta}$-Anosov representations below. (See the paragraph before Fact 2.34 of [24].)
$3.$ $P$-Anosov representations and dominations
Let $\rho:\Gamma_{M}\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ denote a representation. Let $w(g)$ denote the word length of $g$. Suppose that there exists an integer $k$, $1\leq k\leq n-k+1\leq n$, so that the following hold for a constant $A,C>0$:
$\frac{a_{k}(\rho(g))}{a_{k+1}(g)}\geq C\exp(Aw(g))\hbox{ and }\frac{a_{n-k}(\rho(g))}{a_{n-k+1}(g)}\geq C\exp(Aw(g)).$
In this case, we say that $\rho$ is $k$-dominated for $k\leq n/2$ (see Bochi-Potrie-Sambarino [5]).
We will use $\left|\mkern-2.0mu\left|\cdot\right|\mkern-2.0mu\right|$ to indicate the Euclidean norm in a maximal flat in a symmetric space $X$. (See Example 2.12 of [27].) We will use the following notation:
$\begin{split}\vec{a}(g)&:=(a_{1}(g),\dots,a_{n}(g)),\\ \log\vec{a}(g)&:=(\log a_{1}(g),\dots,\log a_{n}(g))\in{\rm Cl}(\mathfrak{a}^{+}_{n}),\hbox{Cartan projection}\\ \log\vec{\lambda}(g)&:=(\log\lambda_{1}(g),\dots,\log\lambda_{n}(g))\in{\rm Cl}(\mathfrak{a}^{+}_{n})\hbox{ Jordan projection}\end{split}$
for the singular values $a_{i}(g)$ and the modulus $\lambda_{i}(g)$ of the eigenvalues of $g$ for $g\in{\mathsf{GL}}(n,{\mathbb{R}})$.
Given a set $A$, we say that two functions $f,g:A\rightarrow{\mathbb{R}}$ are compatible if there exists a uniform constant $C>1$ such that $C^{-1}(x)<f(x)<Cg(x)$ for all $x\in A$.
Lemma 3.1.
Let a finitely generated group $G$ have a representation $\rho:G\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$.
Then the following are equivalent:
  • $\rho$ is $P$-Anosov in the Guéritaud-Guichard-Kassel-Wienhard sense [24] for some parabolic group $P=P_{\theta}$ for an index set $\theta$ containing $k$ and nkn-k.
  • $\rho$ is a $k$-dominated representation for $1\leq k\leq n/2$.
Proof.
This follows by Theorem 1.3(3) of [24] since the set $\theta$ contains $k$ and nkn-k. ∎
4. The proof of Theorem 1.1
4.1. $k$-convexity
We will relate the partial hyperbolic property in the singular-value sense to that in the bundle sense following [5]. We need this modification since we have to have actual expanding and shrinking under the flow.
We continue to assume that $\Gamma_{M}$ is hyperbolic. Let $\rho:\Gamma_{M}\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ be a representation. We say that the representation $\rho$ is $k$-convex if there exist continuous maps $\zeta:\partial_{\infty}\Gamma_{M}\rightarrow\mathcal{G}_{k}({\mathbb{R}}^{n})$ and $\theta:\partial_{\infty}\Gamma_{M}\rightarrow\mathcal{G}_{n-k}({\mathbb{R}}^{n})$ such that the following hold:
(transversality):
for every $x,y\in\partial_{\infty}\Gamma_{M},x\neq y$, we have $\zeta(x)\oplus\theta(y)={\mathbb{R}}^{n}$, and
(equivariance):
for every $\gamma\in\Gamma_{M}$, we have
$\zeta(\gamma x)=\rho(\gamma)\zeta(x),\theta(\gamma x)=\rho(\gamma)\theta(x),x\in\partial_{\infty}\Gamma_{M}.$
Using the representation $\rho$, it is possible to construct a linear flow $\psi_{t}$ over the geodesic flow $\phi_{t}$ of the Gromov flow space $\partial^{(2)}_{\infty}\hat{M}\times{\mathbb{R}}$ as follows. Consider the lifted geodesic flow $\hat{\phi}_{t}$ on $\partial^{(2)}_{\infty}\hat{M}\times{\mathbb{R}}$, and define a linear flow on $\hat{E}:=(\partial^{(2)}_{\infty}\hat{M}\times{\mathbb{R}})\times{\mathbb{R}}^{n}$ by: $\hat{\psi}_{t}(x,\vec{v})=(\phi_{t}(x),\vec{v})$ for $x\in\partial^{(2)}_{\infty}\hat{M}\times{\mathbb{R}}$. Now consider the action of $\Gamma_{M}$ on $\hat{E}$ given by:
$\gamma\cdot\left(x,\vec{v}\right):=\left(\gamma(x),\rho(\gamma)(\vec{v})\right).$
It follows that $\hat{\psi}_{t}$ induces in a flow on $E:={\mathbb{R}}^{n}_{\rho}=\hat{E}/\Gamma_{M}$.
When the representation $\rho$ is $k$-convex, by equivariance, there exists a $\phi_{t}$-invariant splitting of the form $E_{\rho}=Z\oplus\Theta$; it is obtained by taking the quotient of the bundles
$\hat{Z}(x)=\zeta(x_{+})\hbox{ and }\hat{\Theta}(x)=\theta(x_{-})\hbox{ for }x\in\partial^{(2)}_{\infty}\hat{M}\times{\mathbb{R}}$
where $x_{+}$ is the forward endpoint of the complete isometric geodesic through $x$ and $x_{-}$ is the backward endpoint of the complete isometric geodesic through $x$.
We say that a $k$-convex representation is $k$-Anosov in the bundle sense if the splitting $E_{\rho}=Z\oplus\Theta$ is a dominated splitting for the linear bundle automorphism $\psi_{t}$, with $Z$ dominating $\Theta$ in the terminology of [5]. This is equivalent to the fact that the bundle ${\rm Hom}(Z,\Theta)$ is uniformly contracted by the flow induced by $\psi_{t}$.
Suppose that we further require for a $k$-dominated representation $\rho$ for $k\leq n/2$:
  • $a_{p}(\rho(g))\geq C^{-1}\exp(Aw(g))$ for $p\leq k$, and
  • $a_{r}(\rho(g))\leq C\exp(-Aw(g))$ for $r\geq n-k+1$ for some constants $C>1,A>0$.
Then we say that $\rho$ is partially hyperbolic in the singular-value sense with the index $k$.
We obtain by Lemma 3.1:
Lemma 4.1.
If $\rho$ is partially hyperbolic for an index $k$, $k\leq n/2$, in the singular-value sense, then $\rho$ is $P_{\theta}$-Anosov for the index set $\theta$ containing $k$ and nkn-k. ∎
Recall $\hat{M}$ the cover of $M$ with the deck transformation group $\Gamma_{M}$. Recall from Section 2.2 that we can identify $\partial_{\infty}\Gamma_{M}$ for the Cayley graph $\hat{\Gamma}_{M}$ of $\Gamma_{M}$ with $\partial_{\infty}\hat{M}$ by Theorem 11.108 of [14]. Hence, we can identify $\partial_{\infty}^{(2)}\Gamma_{M}$ with $\partial^{(2)}_{\infty}\hat{M}$. The Gromov flow space $\partial_{\infty}^{(2)}\Gamma_{M}\times{\mathbb{R}}$ is identified with $\partial^{(2)}_{\infty}\hat{M}\times{\mathbb{R}}$. By directly following Proposition 4.9 of Bochi-Potrie-Sambarino [5], we obtain:
Theorem 4.2.
Let $\rho:\Gamma_{M}\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ be a representation. Let $\rho$ be partially hyperbolic with an index $k$, $k<n/2$, in the singular-value sense. Then $\rho$ is partially hyperbolic in the bundle sense with the index $k$.
Proof.
By Lemmas 4.1 and 3.1, $\rho$ is $k$-dominated.
Bochi, Potrie, and Sambarino [5] construct subbundles $Z$ and $\Theta$ of $E_{\rho}$ where $Z$ dominates $\Theta$ and $\dim Z=k,\dim\Theta=n-k$.
Now, $\rho$ is also nkn-k-dominated. By Lemma 4.1, we obtain a new splitting of the bundle $E_{\rho}=Z^{\prime}\oplus\Theta^{\prime}$ over $\partial_{\infty}^{(2)}\hat{M}\times{\mathbb{R}}$ where $\dim Z^{\prime}=n-k$ and $\dim\Theta^{\prime}$ is $k$. Furthermore, $Z^{\prime}$ dominates $\Theta^{\prime}$. Then $Z$ is a subbundle of $Z^{\prime}$, and $\Theta^{\prime}$ is a subbundle of $\Theta$ by Proposition 2.1 of [5]. We now form bundles $Z,Z^{\prime}\cap\Theta,\Theta^{\prime}$ over $\partial_{\infty}^{(2)}\hat{M}\times{\mathbb{R}}$. The dominance property (1) of Definition 1.1 follows from those of $Z,\Theta$ and $Z^{\prime},\Theta^{\prime}$.
There are exponential expansions of $a_{i}(g)$ for $1\leq i\leq k$ by the partially hyperbolic condition of the premise. Now (iii)(a) of Definition 1.1 follows: Define $U_{p}(A)$ as the sum of eigenspaces of $\sqrt{AA^{*}}$ corresponding to $p$ largest eigenvalues. Define $S_{n-p}(A):=U_{n-p}(A^{-1})$. Since $\hat{M}/\Gamma_{M}$ is compact, Theorem 2.2 of [5] shows
$U_{k}(\psi^{t}_{\phi^{-1}_{t}(x)})\rightarrow\hat{Z}(x)\hbox{ and }S_{n-k}(\psi^{t}_{x})\rightarrow\hat{\Phi}(x)\hbox{ uniformly for $x\in\hat{\Gamma}_{M}$ as }t\rightarrow\infty.$
Since by the paragraph above Theorem 2.2 of [5], we obtain
[…]
As $t\rightarrow\infty$, $\phi_{t}$ passes the images of the fundamental domain under $\Gamma_{M}$. Let $g_{i}$ denote the sequence of elements of $\Gamma_{M}$ arising in this way. Since $a_{k}(g_{i})$ grows exponentially, the expansion properties of $Z$ follow by the above paragraph. (iii) of Definition 1.1 follows up to reversing the flow.
By the quasi-isometry of $\mathbf{UC}\hat{M}$ with $\partial^{(2)}\hat{M}\times{\mathbb{R}}$ in Theorem 2.6, we construct our pulled-back partially hyperbolic bundles over $\mathbf{UC}\hat{M}$. ∎
4.2. Promoting $P$-Anosov representations to partially hyperbolic ones
This subsection contains the main results of this paper. The complication is that the maximal abelian subgroup of the Zariski closure of the linear part of the holonomy group sometimes embeds not linearly but piecewise linearly. In addition, we have to understand the case where the Zariski closure is not reductive.
[…]
We remark that we put $\mathfrak{a}_{n}$ as the diagonal matrices, and the Jordan projection here is same as the Cartan projection. We will use $\vec{\lambda}$ for $\mathfrak{a}_{n}$ following the convention.
We worked out with the following proposition with the help of Danciger and Stecker:
Proposition 4.3.
Assume the following:
  • $G$ is a finitely generated group.
  • $\rho:G\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ is a $P_{\theta}$-Anosov representation for some index set $\theta=\theta^{*}$ containing $k$ and nkn-k for $1\leq k\leq n/2$.
    • (i)
      $\rho(g)$ for each $g\in G$ has $1$ as the norm of an eigenvalue, or, alternatively,
    • (ii)
      there exists a constant $C>0$ such that $|\log a_{i_{g}}(g)|<C$ for all $g\in\pi_{1}(M)$ where the index $1\geq i_{g}\geq n$ may depend on $g$.
  • The Zariski closure $Z$ of the image of $\rho$ is a connected reductive Lie group.
Then $\rho$ is partially hyperbolic with the index $k$ in the singular-value sense, and $k<n/2$.
Proof.
[…]
Let $B:=\mathcal{BC}_{Z}(\rho(G))$ denote the Benoist cone of $\rho(G)$ in $Z$. Let $S$ be a generating subset of $\rho(G)$. Then $B$ is a span of the joint spectrum $J(S)$ and 0 as stated immediately after Theorem 1.3 of [7]. The image $\iota_{Z}(B)$ is not convex in general.
Let $S\subset\rho(\pi_{1}(N))$ be a generating set. Under the condition (i), each vector $\vec{\lambda}(\rho(g)),g\in\pi_{1}(N)$ has a coordinate with $1$ as a value. By Theorem 1.6 of [7], we have a geometric convergence
$\vec{\lambda}_{Z}(S^{n})\rightarrow J(S)\hbox{ as }n\rightarrow\infty.$
Thus for each $x\in J(S)$, there is a sequence of elements of $x_{n}\in S^{n}$ converging to $x$. Since $B$ is the span of $J(S)$ and 0 in $a^{+}_{Z}$, we find that there must be an index $i$ where $\iota_{Z}(x)_{i}=0$ for every $x\in B$.
Under the condition (ii), for each element $x\in B$, there is a sequence $\{b_{i}\in S\}$ such that
$\lim_{n\rightarrow\infty}\log\vec{a}_{Z}(b_{1}\dots b_{n})=x.$
We find that there must be an index $i$ where $\iota_{Z}(b)=0$ for every $x\in B$.
From now on, the arguments are the same for both conditions. We denote by
$\mathcal{I}(b)=\{i|\log\lambda_{i}(\iota_{Z}(b))=0\}\hbox{ for }b\in B.$
which is a consecutive set under the first premise. Let S([1,,n])S([1,...,n]) denote the collection of subsets of consecutive elements. We define an integral-interval function
$\mathcal{I}:B-\{0\}\rightarrow S([1,\dots,n])$
given by sending $b\in B$ to $\mathcal{I}(b)$. Since the sequences of ordered norms of eigenvalues of a convergent sequence of linear maps converge, we have the following:
(5) $\lim_{i\rightarrow\infty}\mathcal{I}(x_{i})\subset I(x)\hbox{ whenever }x_{i}\rightarrow x,x_{i},x\in B.$
Let $S$ denote a finite set of generators of $\rho(G)$ and their inverses. By Theorems 1.3 and 1.4 of [7], the sequence $\frac{1}{m}\log\vec{\lambda}_{Z}(S^{m})\subset J(S)$ geometrically converges to a compact convex set $J(S)\subset{\rm Cl}(\mathfrak{a}_{Z}^{+})$ as $m\rightarrow\infty$ (see Section 3.1 of [7]). Moreover, the cone spanned by 0 and $J(S)$ is the Benoist cone $B$. This implies that the set of directions of $\log\vec{\lambda}_{Z}(\rho(G))$ is dense in $B$. (See also [4]). Let $S_{n}$ denote the unit sphere in the Lie algebra $\mathfrak{a}_{n}$ of ${\mathsf{GL}}(n,{\mathbb{R}})$ with respect to the standard coordinates. We obtain
[…]
by Theorem 4.2(3) of [24]. Hence, we obtain
(6) $\log\lambda_{k}(b)-\log\lambda_{k+1}(b)\geq C$
for every $b\in B-\{O\}$ going to an element of $\iota_{Z}(B)\cap S_{n}$ for a constant $C>0$.
(II) Our first major step is to prove $\mathcal{I}(B-\{0\})\subset[k+1,n-k]$: Let $\iota_{n}$ denote the opposite involution of $\mathfrak{a}_{n}$. First, suppose that the rank of $Z$ is $\geq 2$, and hence $\dim\mathfrak{a}_{Z}^{+}\geq 2$.
A diagonal subspace
$\Delta_{i,j}\subset\mathfrak{a}_{n}\hbox{ for }i\neq j,i,j=1,\dots,n,$
is a subspace of the maximal abelian algebra $\mathfrak{a}_{n}$ given by $\log\lambda_{i}-\log\lambda_{j}=0$ for some indices $i$ and $j$. The inverse images under $\iota_{Z}$ in ${\rm Cl}(\mathfrak{a}_{Z}^{+})$ of the diagonal subspaces of $\mathfrak{a}$ may meet $B$. We let $\mathcal{I}_{B}$ the maximal collection of indices $(i,j),i<j$, where $\iota_{Z}^{-1}(\Delta_{i,j})$ contains $B$, which may be empty. Let $\mathcal{I}^{\prime}_{B}$ the collection of indices $(k,l)$, $k<l$, of $\Delta_{k,l}$ not in $\mathcal{I}_{B}$. We call the closures of the components of
$B-\bigcup_{(k,l)\in\mathcal{I}^{\prime}_{B}}\iota_{Z}^{-1}(\Delta_{k,l})$
the generic flat domains of $B$. There are finitely many of these to be denoted $B_{1},\dots,B_{m}$ which are convex domains in $B$.
We make notes of two “discrete continuity” properties due to Danciger and Stecker:
  • In the interior of each $B_{i}$, $\mathcal{I}$ is constant since the $\mathcal{I}$-value is never empty and the change in $\mathcal{I}$-values will happen only in the inverse images of some of $\Delta_{k,l}$.
  • For adjacent regions, their images under $\mathcal{I}$ differ by adding or removing some top or bottom consecutive sets.
Heuristically speaking: the worms leave “traces”.
We now aim to prove that the following never occurs –(*):
There exists a pair of elements $b_{1},b_{2}\in B$ with $b_{1}\in B_{p}^{o}$ and $b_{2}\in B_{q}^{o}$ for some $p,q$ where $\mathcal{I}(b_{1})$ has an element $\leq k$ and $\mathcal{I}(b_{2})$ has an element >k>k.
(The idea is that $k,k+1$ both cannot be in a zero set since they differ by a constant due to Anosov properties.)
We use the discrete version of continuity: Suppose that this (*) is true. Since $B$ is convex with $\dim B\geq 2$, there is a chain of adjacent polyhedral images $B_{l_{1}},\dots,B_{l_{m}}$ where $b_{1}\in B_{l_{1}}^{o}$ and $b_{2}\in B_{l_{m}}^{o}$ where $B_{l_{j}}\cap B_{l_{j+1}}$ contains a nonzero point. For adjacent $B_{l_{p}}$ and $B_{l_{p+1}}$, we have by (5)
(7) $\mathcal{I}(B_{l_{p}}^{o})\cup\mathcal{I}(B_{l_{p+1}}^{o})\subset\mathcal{I}(x)\hbox{ for }x\in{\rm Cl}(B_{l_{p}})\cap{\rm Cl}(B_{l_{p+1}})-\{O\}.$
Considering $\mathcal{I}(B_{l_{j}}^{o})$, $j=1,\dots,m$, and the $\mathcal{I}$-values of the intersections of adjacent generic flat domains, we have $\mathcal{I}(b^{\prime})\ni k,k+1$ for some nonzero element $b^{\prime}$ of $B$ by the connectedness of the $\mathcal{I}$-values as we change generic flat domains to adjacent ones according to (7). Hence, we obtain
(8) $\log\lambda_{k}(b^{\prime})=0\hbox{ and }\log\lambda_{k+1}(b^{\prime})=0.$
Since we can assume $\iota_{Z}(b^{\prime})\in S_{n}$ for $b^{\prime}$ in (8) by normalizing, we obtain a contradiction to (6).
Hence, we proved (*), and only one of the following holds:
  • for all $b\in B$ in the interiors of generic flat domains, every element of $\mathcal{I}(b)$ is $\leq k$ or
  • for all $b\in B$ in the interiors of generic flat domains, every element of $\mathcal{I}(b)$ is >k>k.
Note that the opposition involution $\iota_{n}$ sends $\iota_{Z}(B)$ to itself since the inverse map $g\rightarrow g^{-1}$ preserves $Z$ and the set of eigenvalues of $g$ is sent to the set of the eigenvalues of $g^{-1}$. In the first case, by the opposition involution $\iota_{n}$ preserving $G$, $\mathcal{I}(\iota_{n}(b))$ for $b\in B$ contains an element $\geq n-k+1$. This again contradicts the above. Hence, we conclude $\mathcal{I}(b)\subset[k+1,\dots,n]$ for all $b\in B$ in the interiors of generic flat domains. Acting by the opposition involution and using by similar arguments, we obtain $\mathcal{I}(b)\subset[k+1,\dots,n-k]$ for all $b\in B$ in the interiors of generic flat domains.
Furthermore, we can show that $\mathcal{I}(b)\subset[k+1,n-k]$ for all $b\in B-\{0\}$ since otherwise, we have $\mathcal{I}(b^{\prime})\ni k,k+1$ for some $b^{\prime}\neq 0$ by a similar reason to the above applied to lower-dimensional strata. This is a contradiction as before. In addition, the argument shows $k+1,n-k]\neq\emptyset$ and $k<n/2$.
Now we consider the remaining case where $Z$ has rank $1$. Then the maximal abelian group $A_{Z}$ is $1$-dimensional. For each $g\in Z$, there is an index $i(g)$ for which $\log\lambda_{i(g)}(g)=0$. Thus, $\mathfrak{a}_{Z}$ is in the null space of $\log\lambda_{i}$ for some index $i$. Since $g\mapsto g^{-1}$ preserves $\mathfrak{a}_{Z}$, we have $\log\lambda_{n-i+1}=0$ on $\mathfrak{a}_{Z}$. Since ${\rm Cl}(\mathfrak{a}_{Z}^{+})=\mathfrak{a}_{Z}\cap{\rm Cl}(\mathfrak{a}^{+})$, we also have $\log\lambda_{n-i+1}=0$ on $\mathfrak{a}_{Z}^{+}$. Hence,
$\mathcal{I}(B-\{0\})=[i,\dots,n-i+1]\hbox{ for some index }i,0<i\leq n/2.$
Now by (6), we must have $k<i$. This implies $\mathcal{I}(B-\{0\})\subset[k+1,n-k]$.
(III) To complete the proof, we show the partial hyperbolicity property: We let $\log\vec{a}_{Z}|Z:Z\rightarrow\mathfrak{a}_{Z}^{+}$ denote the Cartan projection of $Z$. Recall
$\iota_{Z}\circ\log\vec{a}_{Z}|\rho(G)=\log\vec{a}|\rho(G).$
Since $\rho(G)$ has a connected reductive Zariski closure, the sequence
$\left\{\frac{\log\vec{a}(\rho(g_{l})))}{w(g_{l})}\right\}$
have limit points only in $\iota_{Z}(B)$ by Theorem 1.3 of [7]. By the conclusion of Step (II), for any sequence $\{g_{l}\}$ in $G$,
$\left\{\frac{\min\{|\log a_{i}(\rho(g_{l}))|\}_{i=k+1,\dots n-k}}{w(g_{l})}\right\}\rightarrow 0,$
uniformly in terms of $w(g_{l})$. Hence, this implies that for any small constant $C>0$, there is $N>0$ such that
$a_{i}(\rho(g))\geq\exp(-Cw(g)),g\in G$
for at least one $i\in[k+1,\dots,n-k]$ provided $w(g)>N$. This means that
$a_{i}(\rho(g))\geq A\exp(-Cw(g))$
for at least one $i\in[k+1,\dots,n-k]$ for all $g\in G$ for a constant $A>0$ depending on $C$. By the $P_{\theta}$-Anosov property of $\rho$ and Lemma 3.1, we have
$\frac{a_{k}(\rho(g))}{a_{k+1}(\rho(g))}\geq A^{\prime}\exp(C^{\prime}w(g)),g\in\Gamma$
for some constants $A^{\prime},C^{\prime}>0$. Since
$a_{k+1}(\rho(g))\geq a_{i}(\rho(g)),i=k+1,\cdots,n-k,\hbox{ for every }g\in G,$
it follows that
$a_{k}(\rho(g))\geq A^{\prime\prime}\exp(C^{\prime\prime}w(g))$
for some constants $C^{\prime\prime},A^{\prime\prime}>0$. Taking inverses, we also showed that $\rho^{\prime}$ is partially hyperbolic in the singular-value sense with the index $k$.
Finally, since [k+1,nk][k+1,n-k] is not empty, $k<n/2$. ∎
For a general representation $\phi:\Gamma\rightarrow G$, we define the semisimplification $\phi^{ss}:\Gamma\rightarrow G$ by post-composing $\phi$ with the projection to the Levi factor of the Zariski closure of $\phi(\Gamma)$. (See Section 2.5.4 of [24] for details.) Also, the Zariski closure of a finite-index group is reductive if so is the Zariski closure of the original discrete group. (See Remark 2.38 of [24].)
Proof of Theorem 1.1.
Suppose that $\rho$ is $P$-Anosov in the sense of [24]. Hence, $\Gamma:=\rho(\pi_{1}(N))$ is word-hyperbolic by [24].
First, suppose that the Zariski closure of the image of the linear holonomy group is reductive. By Lemma 3.1, $\rho$ is a $k$-dominated representation of the index $k$. We take a finite-index normal subgroup $\Gamma^{\prime}$ of $\Gamma$. Assume that the Zariski closure $Z$ of $\rho(\Gamma^{\prime})$ is a connected reductive Lie group. By Proposition 4.3, $\rho|\Gamma^{\prime}$ is partially hyperbolic in the singular-value sense with the index $k$, $k<n/2$. So is $\rho$ since $\Gamma$ is a finite-index extension of $\Gamma^{\prime}$. By Theorem 4.2, $\rho$ is a partially hyperbolic representation in the bundle sense with the index $k$.
Now, we drop the reductive condition of the Zariski closure. Mainly, we have to prove the expanding and contracting properties in the subbundles. The semisimplification $\rho^{ss}$ is $P$-Anosov by Proposition 1.8 of [24]. Since conjugations do not change eigenvalues, $\rho^{ss}(g)$ for each $g\in\Gamma$ has an eigenvalue equal to $1$ by Theorem 5.1. (See also the proof of Lemma 2.40 of [24].) By Proposition 4.3 and Theorem 4.2, $\rho^{ss}$ is a partially hyperbolic representation in the bundle sense with the index $k$, $k<n/2$. By Proposition 1.3 of [24], $\rho$ is $P$-Anosov with respect to a continuous Anosov map $\zeta:\partial_{\infty}\Gamma\rightarrow{\mathcal{F}}$ for the flag variety $\mathcal{F}:={\mathsf{GL}}(n,{\mathbb{R}})/P$.
We can conjugate $\rho$ arbitrarily close to $\rho^{ss}$ in ${\rm Hom}(\Gamma,{\mathsf{GL}}(n,{\mathbb{R}}))$ since the orbit of $\rho^{ss}$ is the closed orbit in the closure of the orbit of $\rho$ under the conjugation action. (See Section 2.5.4 of [24]. ) Let $\zeta_{i}:\partial_{\infty}\Gamma\rightarrow{\mathcal{F}}$ denote the Anosov map for $\rho_{i}$ conjugate to $\rho$ converging to $\rho^{ss}$. Theorem 5.13 of [25] generalizes to ${\mathsf{GL}}(n,{\mathbb{R}})$ since we can multiply a representation by a homomorphism $\Gamma\rightarrow{\mathbb{R}}^{+}$ not changing the $P$-Anosov property to ${\mathsf{SL}}_{\pm}(n,{\mathbb{R}})$-representations. (See Section 2.5.3 of [24]). Hence, Theorem 5.13 of [25] implies that $\zeta_{i}$ converges to $\zeta^{ss}:\partial_{\infty}\Gamma\rightarrow{\mathcal{F}}$ the map for $\rho^{ss}$ in the $C^{0}$-topology.
Let ${\mathbb{V}}^{ss+},{\mathbb{V}}^{ss0}$, and ${\mathbb{V}}^{ss-}$ denote the bundles of the partially hyperbolic decomposition of ${\mathbb{R}}^{n}_{\rho^{ss}}$ as in Definition 1.1. Let ${\mathbb{V}}^{+}_{i},{\mathbb{V}}^{0}_{i}$, and ${\mathbb{V}}^{-}_{i}$ denote the bundles of the decomposition of ${\mathbb{R}}^{n}_{\rho_{i}}$ of respective dimensions $k$, n2kn-2k, and $k$ obtainable from $\zeta_{i}$ since $\rho_{i}$ is $k$-dominated. The single product bundle $\mathbf{UC}\hat{M}\times{\mathbb{R}}^{n}$ covers the direct sums of these bundles.
  • Let $\hat{\mathbb{V}}^{ss+},\hat{\mathbb{V}}^{ss0}$, and $\hat{\mathbb{V}}^{ss-}$ denote the cover of ${\mathbb{V}}^{ss+},{\mathbb{V}}^{ss0}$, and ${\mathbb{V}}^{ss-}$.
  • Let $\hat{\mathbb{V}}^{+}_{i},\hat{\mathbb{V}}^{0}_{i}$, and $\hat{\mathbb{V}}^{-}_{i}$ denote the cover of ${\mathbb{V}}^{+}_{i},{\mathbb{V}}^{0}_{i}$, and ${\mathbb{V}}^{-}_{i}$.
Then the above convergence $\zeta_{i}\rightarrow\zeta^{ss}$ means that
(9) $\hat{\mathbb{V}}^{+}_{i}(x)\rightarrow\hat{\mathbb{V}}^{ss+}(x),\hat{\mathbb{V}}^{0}_{i}(x)\rightarrow\hat{\mathbb{V}}^{ss0}(x),\hbox{ and }\hat{\mathbb{V}}^{-}_{i}(x)\rightarrow\hat{\mathbb{V}}^{ss-}(x)\hbox{ as }i\rightarrow\infty$
in the respective Grassmannian spaces pointwise for each $x\in\mathbf{UC}\hat{M}$.
The convergence is also uniform over every compact subset of $\mathbf{UC}\hat{M}$ since these vector spaces depends only on the endpoints of quasi-geodesic though $x$ and $\hat{M}$ is quasi-isometric to $\partial^{2}_{\infty}\Gamma_{M}\times{\mathbb{R}}$.
We may construct a sequence of the fiberwise metrics $\left|\mkern-2.0mu\left|\cdot\right|\mkern-2.0mu\right|_{{\mathbb{R}}^{n}_{\rho_{i}}}$ so that the sequence of the lifted norms of $\left|\mkern-2.0mu\left|\cdot\right|\mkern-2.0mu\right|_{{\mathbb{R}}^{n}_{\rho_{i}}}$ uniformly converging to that of $\left|\mkern-2.0mu\left|\cdot\right|\mkern-2.0mu\right|_{{\mathbb{R}}^{n}_{\rho^{ss}}}$ on $\mathbf{UC}\hat{M}\times{\mathbb{R}}^{n}$ over each compact subset of $\mathbf{UC}\hat{M}$ using a fixed set of a partition of unity subordinate to a covering with trivializations.
The flow $\phi^{ss}$ on ${\mathbb{R}}_{\rho^{ss}}$ satisfies the expansion and contraction properties in Definition 1.1. Note that the lift $\Phi^{ss}$ of $\phi^{ss}$ and that $\Phi_{i}$ of the flow $\phi_{i}$ for ${\mathbb{R}}^{n}_{\rho_{i}}$ in the product bundle $\mathbf{UC}\hat{M}\times{\mathbb{R}}^{n}$ are the same by construction induced from trivial action on the second factor. Since $M$ is compact, there exists $t_{0}$ such that for $t>t_{0}$,
  • $\Phi^{ss}_{t}|{\mathbb{V}}^{ss+}$ is expanding by a factor $c^{+}>1$
  • $\Phi^{ss}_{t}|{\mathbb{V}}^{ss-}$ is contracting by a factor $c^{-}<1$ with respect to the metric $\left|\mkern-2.0mu\left|\cdot\right|\mkern-2.0mu\right|_{{\mathbb{R}}^{n}_{\rho^{ss}}}$.
Hence for sufficiently large $i$, the flow $\Phi_{i}$ on ${\mathbb{R}}^{n}_{\rho_{i}}$ satisfies the corresponding properties as well as the domination properties for $\left|\mkern-2.0mu\left|\cdot\right|\mkern-2.0mu\right|_{{\mathbb{R}}^{n}_{\rho_{i}}}$ by (9). Hence, $\Phi_{i}$ is partially hyperbolic with the index $k$.
For converse, suppose that $\rho$ is partially hyperbolic with the index $k$. Proposition 4.5 of [5] shows that $\rho$ is $k$-dominated. Lemma 3.1 shows that $\rho$ is $P_{\theta}$-Anosov for $k\in\theta$. Also, $k<n/2$ since the neutral bundle exists. ∎
5. Complete affine manifolds
Let $G$ be a Lie group acting transitively and faithfully on a space $X$. A $(G,X)$-structure on a manifold $N$ is a maximal atlas of charts to $X$ so that the transition maps are in $G$. This is equivalent to $N$ having a pair $({\bf dev},h)$ where $\hat{N}$ is a regular cover of $N$ with a deck transformation group $\Gamma_{N}$ such that
  • $h:\Gamma_{N}\rightarrow G$ is a homomorphism, called a holonomy homomorphism, and
  • ${\bf dev}:\hat{N}\rightarrow X$ is an immersion, called a developing map, satisfying
    ${\bf dev}\circ\gamma=h(\gamma)\circ{\bf dev}$
    for each deck transformation $\gamma\in\Gamma_{N}$.
We say that an $m$-manifold NN with a $(G,X)$-structure is complete if ${\bf dev}:\hat{N}\rightarrow X$ is a diffeomorphism. If $N$ is complete, we have a diffeomorphism $N\rightarrow X/h(\Gamma_{N})$, and $h(\Gamma_{N})$ acts properly discontinuously and freely on $X$. Complete $(G,X)$-structures on $N$ are classified by conjugacy classes of representations $\Gamma_{N}\rightarrow G$ with properly discontinuous and free actions on $X$. (See Section 3.4 of [46].)
Note that the completeness and compactness of $N$ have no relation for some $(G,X)$-structures unless $G$ is in the isometry group of $X$ with a Riemannian metric. (The Hopf-Rinow lemma may fail.)
Now, we move on to the affine structures. Let $\mathds{A}^{n}$ be a complete affine space. Let ${\mathbf{Aff}}(\mathds{A}^{n})$ denote the group of affine transformations of $\mathds{A}^{n}$ whose elements are of the form:
$x\mapsto Ax+{\mathbf{v}}$
for a vector ${\mathbf{v}}\in{\mathbb{R}}^{n}$ and $A\in{\mathsf{GL}}(n,{\mathbb{R}})$. Let $\mathcal{L}:{\mathbf{Aff}}(\mathds{A}^{n})\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ denote the map that sends each element of ${\mathbf{Aff}}(\mathds{A}^{n})$ to its linear part in ${\mathsf{GL}}(n,{\mathbb{R}})$.
An affine $n$-manifold is an $n$-manifold with an $({\mathbf{Aff}}(\mathds{A}^{n}),\mathds{A}^{n})$-structure. A complete affine $n$-manifold is an $n$-manifold NN of the form $\mathds{A}^{n}/\Gamma$ for $\Gamma\subset{\mathbf{Aff}}(\mathds{A}^{n})$ acting properly discontinuously and freely on $\mathds{A}^{n}$.
Theorem 5.1 (Hirsch-Kostant-Sullivan [33]).
Let $N$ be a complete affine $n$-manifold. Let $\rho:\pi_{1}(N)\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ be the linear part of the affine holonomy representation $\rho^{\prime}$. Then $g$ has an eigenvalue equal to $1$ for each $g$ in the Zariski closure $Z(\rho(\pi_{1}(N)))$ of $\rho(\pi_{1}(N))$.
Proof.
We define the function $f:Z(\rho(\pi_{1}(N)))\rightarrow{\mathbb{R}}$ by $f(x)=\det(x-{\rm I})$ for each element $x$. Suppose that $f(\rho(g))\neq 0$, $g\in\Gamma_{M}$. Then $\rho(g)(x)+b=x$ has a solution for each $b$ by Cramer’s rule. Hence, $\rho^{\prime}(g)$ has a fixed point, which is a contradiction. We find that $f$ is zero in the Zariski closure, which implies that each element has at least one eigenvalue equal to $1$. ∎
Proof.
Proof of Collary 1.2 By Theorem 5.1, the main premise of Theorem 1.1 is satisfied. Hence, the result follows. ∎